# Model-agnostic Coreset Selection via LLM-based Concept Bottlenecks

### v1

@@ Abstract
Coreset Selection (CS) identifies a subset of training data that achieves model performance comparable to using the entire dataset. Many state-of-the-art CS methods, select coresets using scores whose computation requires training the downstream model on the entire dataset and recording changes in its behavior on samples as it trains (training dynamics). These scores are inefficient to compute and hard to interpret as they do not indicate whether a sample is difficult to learn in general or only for a specific model. Our work addresses these challenges by proposing an interpretable score that gauges a sample’s difficulty using human-understandable textual attributes (concepts) independent of any downstream model. Specifically, we measure the alignment between a sample’s visual features and concept bottlenecks, derived via large language models, by training a linear concept bottleneck layer and compute the sample’s difficulty score using it. We then use this score and a stratified sampling strategy to identify the coreset. Crucially, our score is efficiently computable without training the downstream model on the full dataset even once, leads to high-performing coresets for various downstream models, and is computable even for an unlabeled dataset. Through experiments on CIFAR-10, CIFAR-100, and ImageNet-1K, we show our coresets outperform random subsets, even at high pruning rates, and achieve model performance comparable to or better than coresets found by training dynamics-based methods.
Machine Learning, ICML
@@ 1 Introduction
Machine learning (ML) pipelines are becoming increasingly intensive regarding their data and compute requirements (Touvron et al., 2023; Achiam et al., 2023). While in line with the empirical neural scaling laws (Kaplan et al., 2020; Hestness et al., 2017; Henighan et al., 2020; Rosenfeld et al., 2019) where a model’s performance improves with the increasing model and training data size, these improvements come at an unsustainable cost of compute/energy. Recently, (Sorscher et al., 2022) showed analytically that data pruning can enable models to achieve an exponential reduction in the test error with an increasing dataset size.
@@
Coreset Selection (CS) (Mirzasoleiman et al., 2020; Guo et al., 2022; Paul et al., 2021; Xia et al., 2022; Maharana et al., 2023; Zheng et al., 2022; Choi et al., 2024) aims to improve the efficiency of training large ML models by pruning a large dataset and retaining only a small subset of representative samples. Most CS methods, use metrics to estimate the difficulty/importance of every training sample and then use a sampling strategy to produce a coreset, that leads to similar performance on a downstream model as training it on the entire dataset. Many state-of-the-art (SOTA) CS methods use the downstream model’s training dynamics — the model’s behavior on a sample during training, to generate a difficulty score for each sample. This enables accurate estimation of the data difficulty but requires training the downstream model on the entire dataset at least once, which can be costly when training a large model on a large dataset. While (Coleman et al., 2019) showed that a coreset selected using training dynamics of a mid-sized proxy model (eg., ResNet-18) is effective for a larger downstream model (eg., ResNet-50), training even such a model is not feasible for large datasets. Moreover, these scores are difficult to interpret as they are downstream model-dependent and cannot inform if the sample will be hard or easy to learn for a different downstream model (without training it first). Thus, we aim to answer the following question “How to efficiently estimate a sample’s difficulty in an interpretable and model-agnostic way i.e. without the knowledge of the architecture or training dynamics of the downstream model?”
@@
Figure 1: : Overview of our approach: We start by prompting the language models to generate concept annotation for $N$ class label names in the dataset and select $k$ most discriminative attributes (per class) to form the concept bottleneck, which are passed through a text encoder ($\mathcal{T}_{enc}$) to obtain the bottleneck embedding matrix ($E_{C}$). The visual information of a training sample $x$ extracted via the visual encoder ($\mathcal{V}_{enc}$) is then aligned to the embedding matrix using a linear concept bottleneck layer trained for $T$ epochs. Our difficulty score for $x$ is then computed as the average margin (i.e., the difference between the likelihood of the correct and other classes) over $T$ epochs. Finally, a coreset is selected using stratified sampling which is then used to train an unknown downstream model.
@@
To address this question, we use Concept Bottleneck Models (CBMs) (Koh et al., 2020; Yuksekgonul et al., 2022) which are effective at making ML models more interpretable. CBMs map a model’s input on to a set of human-understandable concepts, referred to as the “bottleneck” and then use them to make a prediction. However, CBMs require concept annotation for every sample in the training data, which can be costly to obtain. Recently (Yang et al., 2023b; Yan et al., 2023) showed that this limitation of CBMs can be overcome by leveraging the advances in Large Language Models (LLMs) and Vision Language Models (VLMs) that can be prompted to generate concept annotation for the training samples without requiring any task-specific fine-tuning. Our Fig. 1 (block 1) shows the prompt we used to obtain the concept bottleneck using an LLM. Once the bottleneck is formed, we use a VLM (such as CLIP (Radford et al., 2021)) to measure the alignment between the visual features and the concept bottleneck (denoted as concept similarity in Fig. 1 (block 2)). A simple linear concept bottleneck layer is then trained to align the visual and concept features to make predictions on the training samples. We use the average margin of a training sample during the training of this bottleneck layer as our concept-based difficulty score. We use a stratified sampling (Zheng et al., 2022) (block 3) using our score to form the coreset which can then be used to train various downstream models Fig. 1 (block 4). Our method speeds up the computation of the coreset by $15$ times compared to approaches based on the downstream model’s training dynamics across three benchmark datasets. Moreover, compared to coresets found by other training dynamics-free approaches our coresets lead to downstream models with significantly better accuracy (see Sec. 5).
@@
We empirically evaluate the effectiveness of our score on three benchmark datasets: CIFAR-10/100 (Krizhevsky et al., 2009) and Imagenet-1K (Deng et al., 2009). Our results show that downstream models trained on our coresets consistently achieve better accuracy than randomly sampled subsets, especially at high data pruning rates, and achieve performance close to SOTA CS methods. We also show that our approach is effective for the label-free CS problem where the dataset is unlabeled and leads to models with superior performance compared to SOTA label-free CS methods on Imagenet. Since our CS method is model agnostic, we show that our coreset leads to high performance regardless of the architecture of the downstream model, unlike other methods which require training the downstream model on the dataset first to identify the coresets for a new model architecture. Lastly, we show that our concept-based difficulty score is aligned with human intuition, allowing us to interpret why examples are easy/hard, independent of the downstream model. Our main contributions are as follows:
@@
  • We propose a concept-based score for CS that gauges sample difficulty in a model-agnostic way without training a downstream model on the full dataset, speeding up CS by $\approx$15x compared to SOTA methods.
@@
  • Our coresets improve model accuracy by $5$% on average over random subsets at high pruning rates, are competitive to SOTA methods, generalize to various architectures, and are computable without labels.
@@
  • We show that using LLM-generated concepts makes our score human-understandable, enabling a generalizable and data-centric solution for CS.
@@ 2 Related Work
Coreset selection (CS): CS improves the efficiency of model training by selecting a subset of influential samples. Various approaches have been proposed to generate such a subset (Guo et al., 2022). A popular approach uses influence functions (Koh & Liang, 2017; Chatterjee & Hadi, 1986; Liu et al., 2021; Schioppa et al., 2022) which measures the influence of a sample by considering the effect of removing it from the model’s training. While effective, these approaches are computationally costly due to their dependence on higher-order derivatives. Another set of approaches selects a subset by either matching the gradients to those computed on the entire dataset (Mirzasoleiman et al., 2020; Killamsetty et al., 2021) or uses training dynamics of a model (Toneva et al., 2018; Pleiss et al., 2020; Lewis & Catlett, 1994; Culotta & McCallum, 2005; Paul et al., 2021) to compute the importance of a sample. However, such approaches require repeated training of the downstream model to produce accurate importance scores. In comparison, our approach avoids using any knowledge of the downstream model for computing the difficulty scores. Finally, using the dataset’s geometric properties via clustering is another popular choice for CS (Sener & Savarese, 2017; Sorscher et al., 2022; Feldman et al., 2020; Feldman & Langberg, 2011; Huang et al., 2019). However, the high computational complexity due to their dependence on pairwise distances between the samples prohibits their use on large datasets.
@@
Concept-based interpretability approaches: Concepts are defined as high-level semantics that refers to the abstract and human-interpretable meanings of the visual data, such as objects, actions, and scenes, as opposed to low-level features like edges or textures (Wu et al., 2016). Concepts have been used in interpretable computer vision to bridge the gap between human understanding and machine perception in various tasks such as image classification. Such interpretability methods can be broadly classified as post-hoc methods (do not impose any model constraints) or by-design methods. Concept Bottleneck Models (CBMs) extend interpretable-by-design approaches by using human-understandable attributes as an intermediate layer for predictions, as used in few-shot learning (Lampert et al., 2013) and attribute learning (Xu et al., 2020; Russakovsky & Fei-Fei, 2012). While interpretable, CBMs reliance on costly annotations and lower accuracy compared to end-to-end models limit their usage. Post-hoc Concept Bottleneck Models (PCBMs) address these issues by incorporating static knowledge bases (e.g., ConceptNet (Speer et al., 2017)) and residual connections to boost accuracy (Yuksekgonul et al., 2022). Recently (Yang et al., 2023b; Yan et al., 2023) incorporated LLMs to identify the concept bottleneck to make classification more explainable. We build on this literature and use CBMs for CS. See App. A for other related work.
@@ 3.1 Coreset selection (CS) problem formulation
Consider a classification task and data distribution $P$. Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote the dataset of $n$ training examples sampled i.i.d. from the distribution $P$ where $x_{i}$ denotes the data and $y_{i}\in\mathcal{Y}$ denotes the label from a set of $N$ classes. CS (Coleman et al., 2019; Zheng et al., 2022) aims to find a subset $\mathcal{S}$ of $\mathcal{D}$ consisting of $m\leq n$ samples such that the models trained on $\mathcal{S}$ achieve performance comparable to models trained on $\mathcal{D}$. Formally, the CS problem is
$\min_{\mathcal{S}:|\mathcal{S}|=m}\mathbb{E}_{(x,y)\sim P}[\ell(x,y|\theta_{\mathcal{S}})]-\mathbb{E}_{(x,y)\sim P}[\ell(x,y|\theta_{\mathcal{D}})],$ (1)
where $\theta_{\mathcal{D}}$ and $\theta_{\mathcal{S}}$ denote the “downstream model” trained on $\mathcal{D}$ and $\mathcal{S}$ (coreset), respectively and $\ell$ is the loss function.
To find this subset $\mathcal{S}$, previous works have proposed scores to gauge a sample’s difficulty for a model, and are later used to form the coreset. Approaches such as max entropy uncertainty sampling (Lewis & Catlett, 1994; Settles, 2012), and least confidence (Culotta & McCallum, 2005) estimate difficulty using the uncertainty of the model’s predictions on a sample. Another set of approaches such as $k$-center greedy (Sener & Savarese, 2017) uses geometric information of the data to filter out redundant samples. Yet, another set of approaches uses information from the training dynamics of the downstream model to estimate the a difficulty score. Scores such as the forgetting score (Toneva et al., 2018) which is computed as the number of times a sample gets misclassified after being correctly classified earlier during model training, the area under the margin (AUM) (Pleiss et al., 2020) which identifies mislabeled/difficult samples, fall in this category. While approaches based on the training dynamics of the downstream model have achieved SOTA results, the requirement of knowledge/training the downstream model or a relatively big proxy model on the entire dataset at least once is inefficient for large datasets/models, even if done only once. This motivates the need of data centric approaches for sample’s difficulty independent of the downstream model.
@@ 3.2 Concept bottleneck models (CBMs)
Recent advances in language model-guided CBMs utilize an LLM to obtain concept bottlenecks which are then used to predict the labels. These works rely on a pre-trained multi-modal models (such as CLIP (Radford et al., 2021)) which consists of a visual encoder $\mathcal{V}_{enc}$ and a text encoder $\mathcal{T}_{enc}$ that can map images and text to a $d$-dimensional representation space. Let $C=\{c_{1},c_{2},\cdots,c_{N_{C}}\}$ be the set of $N_{C}$ concepts (bottleneck) generated via a LLM, we can then construct a bottleneck embedding matrix $E_{C}\in\mathcal{R}^{N_{C}\times d}$ such that each row of the matrix is mapping of the concept $c\in C$ after passing it through textual encoder $\mathcal{T}_{enc}$. Based on this, a CBM (Yang et al., 2023b) produces a prediction $h(x)=f(g(\mathcal{V}_{enc}(x),E_{C}))$ for a sample $x$ where $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N_{C}}$ computes the similarity of the visual features to each concept in the bottleneck and $f:\mathbb{R}^{d}\rightarrow\Delta$ outputs the probability of each class in the label set $\mathcal{Y}$, where $\Delta$ is a $N$ simplex. We discuss details of $f$ and $g$ in Sec. 4.
@@ 4 Methodology
4.1 Generating the concept bottleneck via LLMs
Since obtaining data with concept annotation is costly we use LLMs to generate concept annotation for the samples. However, generating attributes (word-level concepts) for all the samples in the dataset via LLMs is still costly, hence we generate attribute-level concepts only for class label names. This approach was recently shown to be effective at generating the concept bottleneck for interpretable image classification  (Yan et al., 2023; Yang et al., 2023b). In Figure 1 (block 1), we present the prompts provided to the LLMs to extract the concepts for various class label names. The responses of the LLM are then processed to remove formatting errors to obtain the concept sets. Details of our prompt design, robustness check of different prompts, and examples of the generated attributes are mentioned in App. B. Once the per-class concepts are extracted, we select $k$ discriminative concepts per class (concepts that are unique to a class) to form the concept bottleneck. Our final list of $k$ concepts for a class contains its name and $k-1$ attributes generated by the LLM. In Sec. 5.5 we show an ablation study using different methods and LLMs to obtain the concept sets, and the number of concepts $k$.

### v2

# Coreset Selection via LLM-based Concept Bottlenecks

@@ Abstract
Coreset Selection (CS) aims to identify a subset of the training dataset that achieves model performance comparable to using the entire dataset. Many state-of-the-art CS methods select coresets using scores whose computation requires training the downstream model on the entire dataset first and recording changes in the model’s behavior on samples as it trains (training dynamics). These scores are inefficient to compute and hard to interpret, as they do not indicate whether a sample is difficult to learn in general or only for a specific downstream model. Our work addresses these challenges by proposing a score that computes a sample’s difficulty using human-understandable textual attributes (concepts) independent of any downstream model. Specifically, we measure the alignment between a sample’s visual features and concept bottlenecks, derived via large language models, by training a linear concept bottleneck layer and computing the sample’s difficulty score using it. We then use stratified sampling based on this score to generate a coreset of the dataset. Crucially, our score is efficiently computable without training the downstream model on the full dataset even once, leads to high-performing coresets for various downstream models, and is computable even for an unlabeled dataset. Through experiments on CIFAR-10/100, and ImageNet-1K, we show that our coresets outperform random subsets, even at high pruning rates, and achieve model performance comparable to or better than coresets found by training dynamics-based methods.
@@ 1 Introduction
Machine learning (ML) pipelines are increasingly demanding more data and compute touvron2023llama ; achiam2023gpt to achieve improved performance on various tasks. While in line with empirical neural scaling laws kaplan2020scaling ; hestness2017deep ; henighan2020scaling ; rosenfeld2019constructive where a model’s performance improves with increasing model and training data size, these improvements come at an unsustainable cost of compute/energy. However, recently sorscher2022beyond ; li2024datacomp demonstrated that data pruning plays a crucial role in enabling an exponential reduction in test error with increasing dataset size, underscoring the importance of data quality over quantity.
Figure 1: Computation of training dynamics-based scores require training the original or a big proxy model on the full dataset, to identify the coreset, making them computationally expensive.
@@
Coreset Selection (CS) mirzasoleiman2020coresets ; guo2022deepcore ; paul2021deep ; xia2022moderate ; maharana2023d2 ; zheng2022coverage ; choi2024bws is a data pruning technique that improves the efficiency of model training by pruning a large dataset and retaining only a small subset of representative samples. Most CS methods work by first using a score to estimate the difficulty/importance of every training sample and then using a sampling strategy that forms the coreset. Many state-of-the-art (SOTA) CS methods use the downstream model’s training dynamics — changes in the model’s behavior on a sample over epochs during training, to generate an importance score for each sample. While this enables a good estimation of the data importance, it requires training the downstream model on the entire dataset at least once, which can be costly when training a large model on a large dataset (see Fig. 1). While coleman2019selection showed that a coreset selected using training dynamics of a mid-sized proxy model (eg, ResNet-18) is effective for a larger downstream model (eg, ResNet-50), training even such a model may not be feasible for large datasets. Moreover, since these scores are dependent on training dynamics of a particular downstream model, they are hard to interpret as they do not inform us about the sample’s importance for another downstream model (without training it first). Thus, in our work we tackle the following question: “How to efficiently estimate the importance of training samples for CS, in a way that is independent of the downstream model and avoids training that model on the full dataset?
@@
Figure 2: Overview of our approach: We start by prompting an LLM to generate concept annotation for $N$ class label names in the dataset and select $k$ most discriminative attributes (per class) to form the concept bottleneck, which are passed through a text encoder ($\mathcal{T}_{enc}$) to obtain the bottleneck embedding matrix ($E_{C}$). The visual information of a training sample $x$ extracted via the visual encoder ($\mathcal{V}_{enc}$) is then aligned to $E_{C}$ using a linear concept bottleneck layer ($h$) trained for $T$ epochs. Our difficulty score for a sample $x$ is then computed as the average margin (i.e., the difference between the likelihoods of the correct and other classes) over $T$ epochs. Finally, the coreset is selected via stratified sampling and is used to train downstream models.
@@
To address this, we use Concept Bottleneck Models (CBMs) koh2020concept ; yuksekgonul2022post , which work by mapping a model’s input onto a set of human-understandable concepts, referred to as the “bottleneck” and use them to make a prediction. However, CBMs require concept annotation for every sample in the dataset, which can be costly to obtain. Recently yang2023language ; yan2023learning showed that off-the-shelf Large Language Models (LLMs) and Vision Language Models (VLMs) can be prompted to obtain concept annotations for training samples without requiring any task-specific fine-tuning (see Fig. 2 (block 1)). Once the bottleneck is formed, we use a VLM (e.g., CLIP radford2021learning ) to measure the alignment between the visual features and the concept bottleneck (denoted as concept similarity in Fig. 2 (block 2)). A linear concept bottleneck layer is then trained to align the visual and concept features while classifying samples. We then use the average margin of a sample while training the bottleneck layer as our concept-based importance score. Finally, we form the coreset using stratified sampling zheng2022coverage (block 3) based on our score, which can then be used for training downstream models (Fig. 2 (block 4)).
@@
We empirically evaluate the effectiveness of our score on three benchmark datasets: CIFAR-10/100 krizhevsky2009learning and Imagenet-1K deng2009imagenet . Our results show that downstream models trained on our coresets consistently achieve better accuracy than randomly sampled subsets, especially at high data pruning rates, and achieve performance close to SOTA CS methods. We also show that our approach is effective for the label-free CS problem where the dataset is unlabeled and leads to models with superior performance compared to SOTA label-free CS methods, especially on Imagenet. Since our CS method is independent of the downstream model, we show that our coreset leads to high performance regardless of the architecture of the downstream model. Moreover, our method speeds up the computation of the coreset by $\approx 8$ times compared to approaches based on the training dynamics of the downstream model. Lastly, we show that our concept-based difficulty score provides an intuitive explanation of why examples are easy/hard, independent of the downstream model. Our main contributions are summarized below:
@@
  • We propose a concept-based score that efficiently computes a training data sample’s importance for CS without training a downstream model on the full dataset, even once.
@@
  • Our coresets improve accuracy by $\approx 5$% over random subsets at high pruning rates, are competitive to SOTA methods, transfer to various architectures, and are computable without labels.
@@
  • We show that using CBMs with LLM-generated concepts makes our score interpretable, enabling a data-centric solution for identifying coresets.
@@ 2 Related Work
Coreset selection (CS). CS improves the efficiency of model training by selecting a subset of influential samples. Various approaches have been proposed to generate such a subset guo2022deepcore . A popular approach uses influence functions koh2017understanding ; chatterjee1986influential ; liu2021influence ; schioppa2022scaling which measures the influence of a sample by considering the effect of removing it from the model’s training. While effective, these approaches are computationally costly due to their dependence on higher-order derivatives. Approaches that use the dataset’s geometric properties such as are another popular choice for CS sener2017active ; sorscher2022beyond ; feldman2020turning ; feldman2011unified ; huang2019coresets . However, the high computational complexity due to their dependence on pairwise distances between the samples prohibits their use on large datasets. Another set of approaches select a subset by either matching the gradients to those computed on the entire dataset mirzasoleiman2020coresets ; killamsetty2021grad or use training dynamics of a model toneva2018empirical ; pleiss2020identifying ; lewis1994heterogeneous ; culotta2005reducing ; paul2021deep to compute the importance of a sample. However, such approaches require repeated training of the downstream model to produce accurate importance scores. In comparison, our approach avoids using any knowledge of the downstream model for computing the difficulty scores. Unlike our work which focuses on data pruning, a different line of work killamsetty2023milo ; tukan2023provable improves the training efficiency by adaptively selecting the best subset of training data (see App. A).
@@
Concept-based interpretability approaches. Concepts are defined as high-level semantics that refers to the abstract and human-interpretable meanings of the visual data, such as objects, actions, and scenes, as opposed to low-level features like edges or textures wu2016valueexplicithighlevel . Concepts have been used in interpretable computer vision to bridge the gap between human understanding and machine perception in various tasks such as image classification. Such interpretability methods can be broadly classified as post-hoc methods (do not impose any model constraints) or by-design methods (see App. A). Concept Bottleneck Models (CBMs) extend interpretable-by-design approaches by using human-understandable attributes as an intermediate layer for predictions, as used in few-shot learning lampert2013attribute and attribute learning xu2020attribute ; russakovsky2012attribute . While interpretable, CBMs reliance on costly annotations and lower accuracy compared to end-to-end models limit their usage. Post-hoc Concept Bottleneck Models (PCBMs) address these issues by incorporating static knowledge bases (e.g., ConceptNet speer2017conceptnet ) and residual connections to boost accuracy yuksekgonul2022post . Recently yang2023language ; yan2023learning incorporated LLMs to identify the concept bottleneck making classification more explainable. We build on this and use CBMs for CS.
@@ 3.1 Coreset selection (CS) problem formulation
Consider a classification task and data distribution $P$. Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote the dataset of $n$ training examples sampled i.i.d. from the distribution $P$ where $x_{i}$ denotes the data and $y_{i}\in\mathcal{Y}$ denotes the label from a set of $N$ classes. CS coleman2019selection ; zheng2022coverage aims to find a subset $\mathcal{S}$ of $\mathcal{D}$ consisting of $m\leq n$ samples such that the models trained on $\mathcal{S}$ achieve performance comparable to models trained on $\mathcal{D}$. Let $\theta_{\mathcal{D}}$ and $\theta_{\mathcal{S}}$ denote the “downstream model" trained on $\mathcal{D}$ and $\mathcal{S}$ (coreset), respectively and $\ell$ be the loss function then the CS problem is as follows
$\min_{\mathcal{S}:\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=m}\mathbb{E}_{(x,y)\sim P}[\ell(x,y|\theta_{\mathcal{S}})]-\mathbb{E}_{(x,y)\sim P}[\ell(x,y|\theta_{\mathcal{D}})].$ (1)
To find this subset $\mathcal{S}$, previous works have proposed scores that gauge a sample’s difficulty for a model. Approaches such as max entropy uncertainty sampling lewis1994heterogeneous ; Settles_2012 , and least confidence culotta2005reducing estimate difficulty using the uncertainty of the model’s predictions on a sample. Another set of approaches such as $k$-center greedy sener2017active uses geometric information of the data to filter out redundant samples. Yet, another set of approaches uses training dynamics of the downstream model to estimate the difficulty score. Scores such as the forgetting events toneva2018empirical computed as the number of times a sample gets misclassified after being correctly classified earlier during model training, and the area under the margin (AUM) pleiss2020identifying which identifies mislabeled/difficult samples, fall in this category.
@@ 3.2 Concept bottleneck models (CBMs)
Recent advances in language model-guided CBMs utilize an off-the-shelf LLM to obtain concept bottlenecks which are then used to predict the labels. These works rely on pre-trained multi-modal models (such as CLIP radford2021learning ) which consists of a visual encoder $\mathcal{V}_{enc}$ and a text encoder $\mathcal{T}_{enc}$ that can map images and text to a $d$-dimensional representation space. Let $C=\{c_{1},c_{2},\cdots,c_{N_{C}}\}$ be the set of $N_{C}$ concepts (bottleneck) generated via an LLM, we can then construct a bottleneck embedding matrix $E_{C}\in\mathcal{R}^{N_{C}\times d}$ such that each row of the matrix is mapping of the concept $c\in C$ after passing it through textual encoder $\mathcal{T}_{enc}$. Based on this, a CBM yang2023language produces a prediction $h(x)=f(g(\mathcal{V}_{enc}(x);E_{C}))$ for a sample $x$ where $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N_{C}}$ computes the similarity of the visual features to each concept in the bottleneck and $f:\mathbb{R}^{N_{C}}\rightarrow\Delta$ outputs the probability of each class in the label set $\mathcal{Y}$, where $\Delta$ is a $N$ simplex. We discuss details of $f$ and $g$ in Sec. 4.
@@ 4 Methodology
4.1 Generating the concept bottleneck via off-the-shelf LLMs
Since obtaining data with concept annotation is costly, we use LLMs to generate concept annotation for the samples. However, generating attributes (word-level concepts) for all the samples in the dataset via LLMs is still costly, hence we generate word-level concepts only for class label names. This approach was recently shown to be effective at generating the concept bottleneck for interpretable image classification  yan2023learning ; yang2023language . In Figure 2 (block 1), we present the prompts provided to the LLMs to extract the concepts for various class label names. The responses of the LLM are then processed to remove formatting errors to obtain the concept sets. Once the per-class concepts are extracted, we select $k$ discriminative concepts per class (concepts that are unique to a class) to form the concept bottleneck. Our final list of $k$ concepts for a class contains its class name and $k-1$ attributes generated by the LLM. Details of our prompt design, robustness check of different prompts, and examples of the generated attributes are mentioned in App. B.